# When Attention Collapses: How Degenerate Layers in LLMs Enable Smaller, Stronger Models

### v3

@@ F.1 Training details of GPT-2 models
  • Model: 16-layer and 12-layer GPT-2 medium variants
  • Softmax temperature: 1
  • $\alpha$: 0.6
  • Batch size: 50K tokens
  • Learning rate: $3\times 10^{-4}$,
  • Warmup steps: 2K,
  • Scheduler type: cosine decay to $\frac{1}{10}$ of max learning rate,
  • Weight decay: 0.1,
  • Gradient clipping value: 1,
  • Total training steps: 50K
Models Layers Hidden Size Heads Variant
@@
GPT2-large(770M) 36 1280 20 Original
GPT2-large(680M) 32 1280 20 Original
GPT2-medium(355M) 24 1024 16 Original
[…]
GPT2-medium 16 512 8 half width
[…]
GPT2-large 18 1280 20 Ours
GPT2-large 16 1280 20 Ours
GPT2-medium 16 1024 16 Ours
Table 11: Overview of all the GPT2 models used in this study and their architectural configurations. The model configurations of stacking and hybrid stacking are same as our variants.
F.2 Training details of 1.5B OpenLLaMA model
Small base LMs trained with 1B data
We present our main results with Our-1.5B model trained with an existing OpenLLaMA version 1 Geng & Liu (2023) and 1 B tokens randomly sampled from 1T redpajama version1 data. The hyper-parameters related to this model is provided below. It is important to note that our claim that we only use 1 GPU for less than 12 hours to train Our-1.5 B model is specific to models derived using $\mathsf{Inheritune}$ with 1B data. Next we also train multiple sub-models as shown in Figure 20 the training details remains consistent with that of the initial model discussed earlier. However we observe that increasing the number of layers in a sub-model also increase the training time.
Hyper-parameter details of our 1.5B base LM derived using OpenLLaMA-3B as refernce LM:
  • Training tokens: 1B
  • Training epochs: 8
  • Training steps: 64K
  • Learning rate: $3\times 10^{-4}$
  • Scheduler: Cosine
  • Weight decay: 0.1
  • Optimizer: AdamW
  • Warm up steps: 1000
  • Batch size: 131K
  • GPU count: 1
  • GPU type: A6000
  • GPU hours: $\sim$ 8 hours
  • GPU hours/epoch: $\sim$ 54 minutes
Training details of small base LMs with 50B data.
We also trained our 1.5B model with larger subsets of data as shown in Figure 21.It is important to note that all the intermediate tokens until 50B are intermediate checkpoints of a single training run. Some of the key hyper-parameters of our training runs are discussed below. We have also trained three variants of small base LMs utilizing 3 different reference base LMs namely OpenLLaMA-3B, OpenLLaMA-7B and LLaMA2-7B. For target LMs developed with OpenLLaMA-3B we use $n$=13 i.e. 13 layers. For target LMs developed using reference LMs of 7B parameters we use $n$=7 i.e. 7 layers. The training hyper-parameters remains consistent across all the models trained with 50B subset of the pre-train data.
Training hyper-parameters of our target 1.5B and 1.6B small base LMs:
  • Training tokens: 50B
  • Training epochs: $\sim$1
  • Training steps: 191K
  • Learning rate: $3\times 10^{-4}$
  • Scheduler: Cosine
  • Weight decay: 0.1
  • Optimizer: AdamW
  • Warm-up steps: 1000
  • Batch size: 131K tokens
  • GPU count: 1
  • GPU type: A100
  • GPU hours: $\sim$18 hours

### v4

@@ H.1 Causal Attention Rank-1 Collapse and Vanishing Gradients
Definition (attention collapse).
Let $A\in\mathbb{R}^{T\times T}$ be an attention matrix (row-stochastic: $A_{ij}\geq 0$ and $\sum_{j=1}^{T}A_{ij}=1$ for each row $i$). Fix an index $j^{\star}\in\{1,\ldots,T\}$ that is unmasked for all rows (e.g., $j^{\star}=1$ for the BOS token). We say that $A$ is $\varepsilon$-sink-collapsed to $j^{\star}$ if, for every row $i\in\{1,\ldots,T\}$,
$\sum_{j\neq j^{\star}}A_{ij}\;\leq\;\varepsilon.$ (6)
Equivalently, each row places at least $1-\varepsilon$ probability mass on the same column $j^{\star}$, i.e.,
$A_{ij^{\star}}\;=\;1-\sum_{j\neq j^{\star}}A_{ij}\;\geq\;1-\varepsilon,\qquad\forall i\in\{1,\ldots,T\}.$ (7)
Theorem H.1 (Attention collapse to rank-1 sink and vanishing gradients).
Consider a single-head causal self-attention module with $A=\mathrm{softmax}(\frac{1}{\sqrt{d}}QK^{\top}+M)$ and $O=AV$. If $A$ is $\varepsilon$-sink-collapsed to some $j^{\star}$ as in equation 6, then:
(i) Rank-1 collapse. Let $A_{0}\triangleq\mathbf{1}e_{j^{\star}}^{\top}$ (rank-11). Then
$\|A-A_{0}\|_{F}\leq\varepsilon\sqrt{2T}\;\Rightarrow\;\sigma_{2}(A)\leq\|A-A_{0}\|_{2}\leq\|A-A_{0}\|_{F}\leq\varepsilon\sqrt{2T},$ (8)
so $A$ is numerically rank-$1$ where $\varepsilon>0$ is small.
(ii) Vanishing gradients through softmax. For each row $i$, letting $a^{(i)}$ be the entries of that row (restricted to unmasked indices),
$\frac{\partial\mathcal{L}}{\partial\tilde{a}^{(i)}}=\Bigl(\mathrm{diag}(a^{(i)})-a^{(i)}{a^{(i)}}^{\top}\Bigr)\;\frac{\partial\mathcal{L}}{\partial a^{(i)}},\quad\Bigl\|\frac{\partial\mathcal{L}}{\partial\tilde{a}^{(i)}}\Bigr\|_{2}\leq 2\varepsilon\;\Bigl\|\frac{\partial\mathcal{L}}{\partial a^{(i)}}\Bigr\|_{2}.$ (9)
Thus, once rows are near one-hot ($\varepsilon\!\to\!0$), the gradient to logits vanishes.
(iii) Vanishing gradients to $W_{Q}$ and $W_{K}$. Moreover,
$\Bigl\|\frac{\partial\mathcal{L}}{\partial W_{Q}}\Bigr\|_{F}\;\leq\;\frac{2\varepsilon}{\sqrt{d}}\;\|X\|_{2}\;\|K\|_{2}\;\Bigl\|\frac{\partial\mathcal{L}}{\partial A}\Bigr\|_{F},\quad\Bigl\|\frac{\partial\mathcal{L}}{\partial W_{K}}\Bigr\|_{F}\;\leq\;\frac{2\varepsilon}{\sqrt{d}}\;\|X\|_{2}\;\|Q\|_{2}\;\Bigl\|\frac{\partial\mathcal{L}}{\partial A}\Bigr\|_{F},$ (10)
so parameter updates to queries/keys shrink linearly with the collapse level $\varepsilon$.
Proof.
We prove the three claims sequentially.
(i) Let $A_{0}\;=\;\mathbf{1}e_{j^{\star}}^{\top}\in\mathbb{R}^{T\times T}$. Clearly $A_{0}$ has rank $1$. As the Frobenius norm is the $\ell_{2}$ norm aggregated over rows, it follows that $\|A-A_{0}\|_{F}^{2}\;=\;\sum_{i=1}^{T}\bigl\|a^{(i)}-e_{j^{\star}}\bigr\|_{2}^{2}$. Fix a row $a\in\mathbb{R}^{T}$ (without loss of generality we omit the superscript $i$), under $\varepsilon$-sink-collapse, $\sum_{j\neq j^{\star}}a_{j}\leq\varepsilon$. Since $A$ is an attention matrix (row-stochastic), then $\sum_{j=1}^{T}a_{j}=1$ and $a_{j}\geq 0$, hence $a_{j^{\star}}\;=\;1-\sum_{j\neq j^{\star}}a_{j}\;\geq\;1-\varepsilon.$ Now consider the squared error $\|a-e_{j^{\star}}\|_{2}^{2}\;=\;(a_{j^{\star}}-1)^{2}+\sum_{j\neq j^{\star}}a_{j}^{2}.$ Here, $(a_{j^{\star}}-1)^{2}\;\leq\;\varepsilon^{2}$ (follows from $a_{j^{\star}}\geq 1-\varepsilon$), and $\sum_{j\neq j^{\star}}a_{j}^{2}\;\leq\;\Bigl(\sum_{j\neq j^{\star}}a_{j}\Bigr)^{2}\;\leq\;\varepsilon^{2}$. Thus, the squared error $\|a-e_{j^{\star}}\|_{2}^{2}\leq 2\varepsilon^{2}$. Summing over $T$ rows, gives
@@
$\text{Hence,}\;\sigma_{2}(A)\leq\|A-A_{0}\|_{2}\leq\|A-A_{0}\|_{F}\leq\varepsilon\sqrt{2T}\quad\text{(following \emph{Weyl's inequality})}.$
(ii) (Gradients are taken only over unmasked logits; masked positions have zero gradient.) Let $g\triangleq\partial\mathcal{L}/\partial a^{(i)}$ and $\tilde{g}\triangleq\partial\mathcal{L}/\partial\tilde{a}^{(i)}$. For the row-wise softmax, the Jacobian is $J=\mathrm{diag}(a)-aa^{\top}$, hence $\tilde{g}=Jg$. Since $J\succeq 0$, $\|J\|_{2}\leq\mathrm{tr}(J)$. Moreover, $\mathrm{tr}(J)=\sum\limits_{i=1}^{T}a_{i}(1-a_{i})=1-\|a\|_{2}^{2}$. Under $\varepsilon$-sink collapse, $a_{j^{\star}}\geq 1-\varepsilon$, so $\|a\|_{2}^{2}\geq(1-\varepsilon)^{2}$ and therefore
$\|J\|_{2}\leq 1-(1-\varepsilon)^{2}\leq 2\varepsilon.$
Thus $\|\tilde{g}\|_{2}\leq\|J\|_{2}\|g\|_{2}\leq 2\varepsilon\|g\|_{2}$, proving equation 9.
(iii) We focus on $W_{Q}$; the argument for $W_{K}$ is analogous. By the chain rule, $\frac{\partial\mathcal{L}}{\partial W_{Q}}=X^{\top}\frac{\partial\mathcal{L}}{\partial Q}$. Let $\bar{G}\triangleq\frac{\partial\mathcal{L}}{\partial\tilde{A}}$. Since $\tilde{A}=\frac{1}{\sqrt{d}}QK^{\top}+M$, we have $\frac{\partial\mathcal{L}}{\partial Q}=\frac{1}{\sqrt{d}}\bar{G}K$. Bounding norms and using $\|X^{\top}H\|_{F}\leq\|X\|_{2}\|H\|_{F}$, (by definition)
$\Bigl\|\frac{\partial\mathcal{L}}{\partial W_{Q}}\Bigr\|_{F}\leq\|X\|_{2}\Bigl\|\frac{\partial\mathcal{L}}{\partial Q}\Bigr\|_{F}\leq\frac{1}{\sqrt{d}}\|X\|_{2}\|\bar{G}\|_{F}\|K\|_{2}.$
From (ii), row-wise $\|\partial\mathcal{L}/\partial\tilde{a}^{(i)}\|_{2}\leq 2\varepsilon\|\partial\mathcal{L}/\partial a^{(i)}\|_{2}$, and therefore $\|\bar{G}\|_{F}\leq 2\varepsilon\|\frac{\partial\mathcal{L}}{\partial A}\|_{F}$. Substituting gives equation 10. ∎
Interpretation and training-time implication.
Theorem H.1 connects rank-1 causal attention collapse (sink behavior) and vanishing gradients: once a head becomes near-deterministic, the softmax Jacobian saturates and the head receives negligible learning signal to recover. Consequently, collapse becomes sticky (hard to recover from) after entering the collapsed regime, subsequent gradient updates tend to be too small to move the head back to a higher rank (non-collapsed) state. Over training, this stickiness can accumulate across layers especially in deeper layers which are more prone to vanishing gradients leading to increasingly persistent rank collapsed attention patterns and progressively smaller query/key weight updates.